# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ 4.3 Tools and techniques to simulate PS-PIDE models
We note that we can obtain the PS-PIDE (7) by taking $A\equiv 0$, the PS-PIDE (17) with $\Pi(\rho):=\rho$ by taking $D\equiv 0$ and $A:=-\mu(y)\partial_{x}\rho$, and the PS-PIDE (20) by taking $D\equiv 0$ and $A:=\chi(y)\partial_{x}S$.
The MOL has been the most popular method for simulating PS-PIDE models, and involves a discretisation of the PS-PIDE in both phenotype and physical space to obtain a high dimensional system of ODEs. The solution of this system is tracked over a computational mesh and integrated in time through an appropriate method.
@@
Naturally, numerical solutions based on the MOL require finite domains [^17] If one needs to simulate a PS-PIDE on an unbounded domain, some alternative approach (such as a pseudospectral method, e.g. Fornberg 102) would need to be invoked., i.e. $\mathcal{X}\subset\mathbb{R}$ and $\mathcal{Y}\subset\mathbb{R}$ in (69). Consider a uniform discretisation of the domain $\mathcal{X}\times\mathcal{Y}$, comprising of $Nx\times Ny$ grid cells each of area $\Delta x\times\Delta y$ and with cell centres $\{x_{i}\}_{i=1,...,Nx}\times\{y_{j}\}_{j=1,...,Ny}$, where $x_{i}=i\Delta x-\frac{1}{2}\Delta x$ ($i=1,...,Nx$) and $y_{j}=j\Delta y-\frac{1}{2}\Delta y$ ($j=1,...,Ny$). We then consider a numerical approximation for the average of $n$, $S$, and $\rho$ across each cell, and let
$n(t,x_{i},y_{j})\approx n_{i,j}(t)\,,\quad\rho(t,x_{i})\approx\rho_{i}(t):=\Delta y\sum_{j=1}^{Ny}n_{i,j}(t)\,,\quad S(t,x_{i})\approx S_{i}(t)\,,$
@@
$M_{i,j}(t):=\,\frac{1}{\Delta x}\left[F^{*}_{i+\frac{1}{2},j}(t)-F^{*}_{i-\frac{1}{2},j}(t)\right],\qquad i=2,...,Nx-1,\;\;j=1,...,Ny,$
@@
$C_{i,j}(t):=\bar{D}\,\frac{n_{i,j-1}(t)-2n_{i,j}(t)+n_{i,j+1}(t)}{(\Delta y)^{2}},\quad i=1,...,Nx,\;\;j=2,...,Ny-1,$
@@
Following the discretisation in phenotype and physical space, one needs to solve the system of $Nx\times Ny$ ODEs (70) by integrating in time. Consider a uniform discretisation of the time interval $[0,T]$ with time-step of size $\Delta t$, i.e. $t_{k}=k\Delta t$ ($k=0,...,Nt$). Introducing the notation $n^{k}_{i,j}:=n_{i,j}(t_{k})$ – and similarly for the other time-dependent variables – a first-order forward difference scheme yields
@@
This is the simplest explicit scheme that one may adopt, being first-order in time; higher-order schemes may be more appropriate if the problem requires higher accuracy, e.g. when observing features as in [172]. Nonetheless, stiffer problems may require an excessively small value for $\Delta t$ to ensure stability in an explicit scheme, e.g. due to a severe CFL constraint [134]. If so, an implicit scheme may be preferable and, as an example, a first-order backward difference for the time derivative yields
@@
The use of implicit solvers is also particularly useful for preserving non-negativity of solutions: crucial, given the usual biological meaning of $n$ and – generally – the structural properties of the PS-PIDE . Explicit solvers may indeed fail at this, for example if the negative part of $R$ becomes particularly large. Implicit-explicit approaches can also be adopted to circumvent this issue, for instance by splitting the reaction term to treat the negative part of $nR$ implicitly and the positive part explicitly [156, 167]. Time-splitting schemes relying on exploiting the WKB ansatz and solving the equation for $u$ instead, in the non-conservative part of the problem, have also been proposed [173].
@@ 5 Challenges and perspectives
PS-PIDEs form a powerful modelling tool to account for the complex phenotypic heterogeneity inherent to a population. With their greater adoption in mathematical models of collective cell migration, various challenges have emerged with respect to modelling extensions, analytical results, numerical methods, and the capacity to connect models with experimental data. We conclude this review with a few perspectives.
  • I. The challenge of cell signalling. PS-PIDE models of cell population dynamics often invoke a somewhat hazy definition of phenotype: a generic variable that defines cell behaviour in a phenomenological manner. In practice, phenotype is linked to the internal signalling state and modulated by environmental factors (e.g. extracellular ligands). ABMs can account for this complexity – e.g. equipping each agent with a detailed ODE model of signalling with the output altering the phenotypic state – and various open source modelling toolkits have been developed with such capabilities (see [192] for an overview). Can similar detail be absorbed within continuous models? As described, environmental heterogeneity can be accounted for in a PS-PIDE model, by variables for the external factors that regulate growth, movement, and phenotype changes (e.g. Ayati et al. 19, Dyson et al. 80, Fiandaca et al. 93, Lorenzi et al. 176, Villa et al. 237, Fiandaca et al. 92, Lorenzi and Painter 172, Lorenzi et al. 168). However, these usually take a black box approach for the intracellular processes that link the extracellular to cell behaviour. More direct definitions for the variability offer a more explicit connection, e.g. associating heterogeneity with tumbling bias in PS-PIDE models for E.coli migration [182, 207]. Fundamental approaches to link structure to signalling could exploit ABM to population-level formulations, e.g. the framework described in Section 4.1; a step here has been taken in [99], where a PS-PIDE model is derived in which phenotype is associated to bound receptor levels, which in turn evolve through the modelling of receptor-ligand binding. PS-PIDE models have been formulated where structuring is according to intracellular signalling pathways connected with EMT [^18] Epithelial-to-mesenchymal transition, where a cells transitions from an epithelial (e.g. high cell-cell adhesion, low motility) to mesenchymal (e.g. cell-extracellular matrix interactions, high motility) state. processes [112], although these do not include spatial movement. Others have addressed a similar challenge with different frameworks, for example by starting from a ‘mesoscopic’ kinetic equation that includes dependency on an internal state variable and employing scaling methods to obtain macroscopic PDE models (but not PS-PIDEs) with terms that depend on the internal signalling state (e.g. Erban and Othmer 88, Engwer et al. 85). Nevertheless, this remains an area for significant work, and will become increasingly necessary if this biological detail is required.
@@
  • III. Extension of other biological movement models. In Section 3 we have illustrated the extension of a number of common PDE models of biological movement into the PS-PIDE form, specifically diffusion-based, pressure-based, and taxis-based models described therein. There are, naturally, further models used to describe different forms of biological movement. Non-local aggregation models [206], wherein advection through physical space is according to a non-local evaluation of surrounding population density, have become a popular tool to describe the impact of direct cell to cell (or animal to animal) interactions – such as cell-cell adhesion – on movement. Variation in adhesion is a common and potent form of phenotypic variation: downregulated cell-cell adhesion can lead to more invasive phenotypes within cancer invasion [244]; differential adhesion is a powerful driver of cellular sorting during tissue patterning [227]. While non-local aggregation models have considered cellular heterogeneity in a discrete manner (e.g. two populations with distinct adhesion, see [206] and references therein), extensions to continuous heterogeneity are natural; steps in this direction have been made, for example see [90]. The resulting equations, though, present a formidable mathematical and numerical challenge: for example, a doubly non-local structure, with integral terms for summation across both space and phenotype.
@@
  • IV. Discrete, continuous, or both? As noted several times, a natural application of PS-PIDEs would be to describe follower-leader type behaviour, for example as can be found in invading cancer cells [244], neural crest migration [190], and wound healing [245]. While a division into followers and leaders is conceptually straightforward, this can be overly simplistic and misleading [228]. A broad spectrum of states may exist between ‘follower’ and ‘leader’, with transitions in between. PS-PIDE models, therefore, can form a natural framework to describe this complexity. But questions then arise regarding the appropriateness of using continuous densities: within such systems, certain states may be confined to small subpopulations – for example, an invasion process driven by relatively few cells at the front. Turning to a fully discrete representation is one possibility, but this may also be suboptimal if other states contain vast numbers of cells. How can we form a bridge, spanning discrete to continuous populations according to phenotypic state? Some hybrid approaches offer tantalising directions, e.g. allowing individual cells to emerge from or merge with some continuous distribution, in a way that ensures mass is conserved [49].
@@
  • V. Self-organisation within phenotype-structured populations. PS-PIDEs are sophisticated equations and present a formidable mathematical challenge. However, in-roads are being made: for example, we illustrated the growing arsenal of techniques to understand phenotype-structured travelling waves. Travelling waves are one well-known dynamic that can arise from PDEs, another being self-organising phenomena. [232] laid the essential groundwork here, via his well known model for morphogenesis. The now standard stability analysis relied on a near homogeneous initial state but, as noted earlier, populations are rarely homogeneous and one could expect phenotypic heterogeneity across even a superficially similar population – e.g. varying levels of signal activity. Extending pattern formation analyses to models that contain phenotype structuring is relatively straightforward under simpler binary phenotypes (e.g. see Painter 197, Pham et al. 201, Macfarlane et al. 185), as the resultant PDEs are of low order and the same standard Turing-type analysis can be applied. But extending these analyses to the non-local PS-PIDE framework has, to today, received little attention.
@@
  • VI. Analysis of qualitative and quantitative properties of solutions. The analysis of PS-PIDEs modelling spatial spread and evolutionary dynamics of forms (7), (17), and (20) is a relatively young research field, wherein a number of fascinating and challenging open problems have emerged. Just to mention a few problems related to aspects we have focused on in this review, neither explicit nor semi-explicit solutions are known, with the exception of the case discussed in Section 4.2.3. This limits our current level of information on the quantitative properties of the solutions to these models. Moreover, formal asymptotic results on concentration phenomena across travelling waves, like those discussed here, have still to be established rigorously. In this we anticipate major difficulties to be posed due to the lack of compactness [186, 139]. Furthermore, there are no rigorous asymptotic results, of the type presented in [70], on the derivation of free-boundary problems of Hele-Shaw type from PS-PIDE models of form (17), when the mobility parameter is a function of the phenotypic state. This problem appears to be far from being closed given that, as of today, related problems remain open even for systems of PDEs with cross-diffusion terms corresponding to the case where the phenotypic state is binary [166, 72]. Solving these and related problems will entail harnessing a range of tools and techniques from across different research areas for the analysis of non-linear and non-local PDEs, thus promoting cross-fertilisation of these areas. Further, it may lead to the development of new mathematical methods that could also be transferable to cognate research fields.
@@
  • VII. PS-PIDEs over networks. Migration not only allows cells to infiltrate neighbouring areas but also disperse to distant sites, as happens in the metastatic spread of cancer. The question of whether cancerous cells that originate in one organ can migrate to, adapt to, and persist within another organ is of manifest interest; related questions arise in an ecological context. In such instances, continuous descriptions of space may no longer be appropriate – the distance between two organs is not easily defined – and formulations on a network may be more appropriate. Steps in this direction have been made, by describing the dynamics of the phenotype density in each network node with a non-local PDE of form (10), and adding sink/source terms for the migration across adjacent edges. Assuming fitness functions in the form of (11)-(12), analytical results extend those outlined in Section 4.2.2 and highlight that spatial movement — even in discrete space settings — may provide the substrate for polymorphism, both under constant [184, 183] or phenotype-dependent migration rates [210]. The long-term persistence (or extinction) of phenotype-structured populations in such modelling framework was investigated in [128, 10], in the context of host-pathogen interactions. Nevertheless, PS-PIDEs over networks currently neglect explicit spatial dynamics within each node – with the exception of a graph-like description of the habitat in each node [42] – and the inclusion of continuous spatial structures for node-specific dynamics is an open research avenue.
@@
  • VIII. Development and analysis of numerical schemes. Relying on structuring over both phenotype and physical space, PS-PIDE models necessitate discretisation over multiple dimensions to be simulated. This aspect, coupled with the presence of non-local and highly non-linear terms, often makes solving numerically the equations comprised in these models prohibitively expensive, especially when high dimensional phenotypic and physical domains are simultaneously considered. As such, a significant challenge is posed when it comes to developing efficient algorithms implementing accurate numerical schemes. As discussed in Section 4.3, approaches to date have typically invoked semi-classical ideas (method of lines, finite difference, finite volume schemes, etc), which benefit from a substantial literature and flexibility. However, whether other approaches – such as pseudospectral methods [102], discontinuous Galerkin methods [57], particle methods [8], and deep-learning algorithms [22] – could be used to generate more efficient algorithms is an intriguing possibility. Beyond the development of efficient algorithms, systematic numerical analysis of discretisation schemes for PS-PIDE models of cell movement of the type considered here is generally lacking in the extant literature, which would provide valuable information on accuracy of numerical solutions; this is certainly another avenue for future research.
@@
  • IX. Bridging the gap between models and data. Recent technological advances have made it possible to access a vast amount of data on the phenotypic properties of cells, both at the single-cell level and across space (see Figure 2(c)), that could be used to validate PS-PIDE models of collective cell migration. While some works already leveraged proteomics data from bulk measurements, of the type shown in panel (ii) of Figure 2(c), for parameter estimation in spatially homogeneous models of phenotypically-structured cell population dynamics [6, 48], this challenge is yet to be tackled in PS-PIDEs models of cell movement. Indeed spatial data come at varying levels of resolution and may be more prone to stochastic variation compared to bulk measurement data, making the task of PDE calibration more involved. This raises the question of practical identifiability of parameters in PS-PIDE models, which is bound to merge complexities found when addressing this in spatially-structured phenotypically homogeneous populations [175] and in well-mixed phenotype-structured populations [30].
@@
Ethics approval Not applicable.
@@
Code availability Code is available upon request.

### v2

@@ 4.3 Tools and techniques to simulate PS-PDE models
We note that from this PS-PDE model we can obtain: the PS-PDE model (7) by taking $A\equiv 0$ and $R(y,\rho,S)\equiv R(y,\rho)$; the PS-PDE model (17) with $\Pi(\rho):=\rho$ by taking $D\equiv 0$, $A(y,\rho,S)\equiv A(y,\rho):=-\mu(y)\partial_{x}\rho$, and again $R(y,\rho,S)\equiv R(y,\rho)$; and the PS-PDE model (20) by taking $D(y)\equiv D$ and $A(y,\rho,S)\equiv A(y,S):=\chi(y)\partial_{x}S$. Python code to solve the PS-PDE model (7) using the scheme described below is available in Open Access (see ‘Code Availability’ at the end of the manuscript).
The MOL has been the most popular method for simulating PS-PDE models, and involves a discretisation of the PS-PDE in both phenotype and physical space to obtain a high-dimensional system of ODEs. The solution of this system is tracked over a computational mesh and integrated in time through an appropriate method.
@@
Naturally, numerical solutions based on the MOL require bounded domains [^19] If one needs to simulate a PS-PDE on an unbounded domain, some alternative approach – such as a pseudospectral method, see e.g. [104] – would need to be invoked., i.e. $\mathcal{X}\subset\mathbb{R}$ and $\mathcal{Y}\subset\mathbb{R}$ in (69). Consider a uniform discretisation of the domain $\mathcal{X}\times\mathcal{Y}$, comprising of $N_{x}\times N_{y}$ grid cells each of area $\Delta x\times\Delta y$ and with cell centres $\{x_{i}\}_{i=1,...,N_{x}}\times\{y_{j}\}_{j=1,...,N_{y}}$, where $x_{i}=i\Delta x-\frac{1}{2}\Delta x$ ($i=1,...,N_{x}$) and $y_{j}=j\Delta y-\frac{1}{2}\Delta y$ ($j=1,...,N_{y}$). We then consider a numerical approximation for the average of $n$, $S$, and $\rho$ across each cell, and let
$n(t,x_{i},y_{j})\approx n_{i,j}(t)\,,\quad\rho(t,x_{i})\approx\rho_{i}(t):=\Delta y\sum_{j=1}^{N_{y}}n_{i,j}(t)\,,\quad S(t,x_{i})\approx S_{i}(t)\,,$
@@
$M_{i,j}(t):=\,\frac{1}{\Delta x}\left[F^{*}_{i+\frac{1}{2},j}(t)-F^{*}_{i-\frac{1}{2},j}(t)\right],\qquad i=2,...,N_{x-1},\;\;j=1,...,N_{y},$
@@
$C_{i,j}(t):=\bar{D}\,\frac{n_{i,j-1}(t)-2n_{i,j}(t)+n_{i,j+1}(t)}{(\Delta y)^{2}},\quad i=1,...,N_{x},\;\;j=2,...,N_{y-1},$
@@
Following the discretisation in phenotype and physical space, one needs to solve the system of $N_{x}\times N_{y}$ ODEs (70) by integrating in time. Consider a uniform discretisation of the time interval $[0,T]$ with time-step of size $\Delta t$, i.e. $t_{k}=k\Delta t$ ($k=0,...,N_{t}$). Introducing the notation $n^{k}_{i,j}:=n_{i,j}(t_{k})$ – and similarly for the other time-dependent variables – a first-order forward difference scheme yields
@@
This is the simplest explicit scheme that one may adopt, being first-order in time; higher-order schemes may be more appropriate if the problem requires higher accuracy, e.g. when investigating features as those considered in [165]. Nonetheless, stiffer problems may require an excessively small value for $\Delta t$ to ensure stability in an explicit scheme, e.g. due to a severe CFL (Courant–Friedrichs–Lewy) constraint [136]. If so, an implicit scheme may be preferable and, as an example, a first-order backward difference for the time derivative yields
@@
The use of implicit solvers is also particularly useful for preserving non-negativity of solutions: crucial, given the usual biological meaning of $n$ and – generally – the structural properties of the PS-PDE . Explicit solvers may indeed fail at this, for example if the negative part of $R$ becomes particularly large. Implicit-explicit approaches can also be adopted to circumvent this issue, for instance by splitting the reaction term to treat the negative part of $nR$ implicitly and the positive part explicitly [167, 174]. Time-splitting schemes relying on exploiting the WKB ansatz and solving the equation for $u$ instead, in the non-conservative part of the problem, have also been proposed [171].
@@ 5 Challenges and perspectives
PS-PDEs form a powerful modelling tool to account for the complex phenotypic heterogeneity inherent to a population. With their greater adoption in mathematical models of collective cell migration, various challenges have emerged with respect to modelling extensions, analytical results, numerical methods, and the capacity to connect models with experimental data. We conclude this review with a few perspectives.
  • I. The challenge of cell signalling. PS-PDE models of cell population dynamics often invoke a somewhat hazy definition of phenotype: a generic variable that defines cell behaviour in a phenomenological manner. In practice, phenotype is linked to the internal signalling state and modulated by environmental factors (e.g. extracellular ligands). ABMs can account for this complexity – e.g. equipping each agent with a detailed ODE model of signalling with the output altering the phenotypic state – and various open source modelling toolkits have been developed with such capabilities – see [189] for an overview. Can similar detail be absorbed within continuous models? As described, environmental heterogeneity can be accounted for in a PS-PDE model, by variables for the external factors that regulate growth, movement, and phenotype changes (e.g. Ayati et al 20, Dyson et al 84, Fiandaca et al 94, Lorenzi et al 170, Villa et al 245, Fiandaca et al 95, Lorenzi and Painter 165, Lorenzi et al 173). However, these usually take a black box approach for the intracellular processes that link the extracellular to cell behaviour. More direct definitions for the variability offer a more explicit connection, e.g. associating heterogeneity with tumbling bias in PS-PDE models for E.coli migration [184, 212]. Fundamental approaches to link structure to signalling could exploit ABM to population-level formulations, e.g. the framework described in Section 4.1; a step here has been taken by [109], who derived a PS-PDE model wherein phenotype is associated to bound receptor levels, which in turn evolve through the modelling of receptor-ligand binding. PS-PDE models have been formulated where structuring is according to intracellular signalling pathways connected with EMT [^20] Epithelial-to-mesenchymal transition, where a cells transitions from an epithelial (e.g. high cell-cell adhesion, low motility) to mesenchymal (e.g. cell-extracellular matrix interactions, high motility) state. processes [118], although these do not include spatial movement. Others have addressed a similar challenge with different frameworks, for example by starting from a ‘mesoscopic’ kinetic equation that includes dependency on an internal state variable and employing scaling methods to obtain macroscopic PDE models (but not PS-PDEs) with terms that depend on the internal signalling state (e.g. Erban and Othmer 89, Engwer et al 87). Nevertheless, this remains an area for significant work, and will become increasingly necessary if this biological detail is required.
@@
  • III. Extension of other biological movement models. In Section 3 we have illustrated the extension of a number of common PDE models of biological movement into the PS-PDE form, specifically diffusion-based, pressure-based, and taxis-based models described therein. There are, naturally, further models used to describe different forms of biological movement. Non-local aggregation models [203], wherein advection through physical space is according to a non-local evaluation of surrounding population density, have become a popular tool to describe the impact of direct interactions – such as cell-cell adhesion – on movement. Variation in adhesion is a common and potent form of phenotypic variation: downregulated cell-cell adhesion can lead to more invasive phenotypes within cancer invasion [243]; differential adhesion is a powerful driver of cellular sorting during tissue patterning [236]. While non-local aggregation models have considered cellular heterogeneity in a discrete manner (e.g. two populations with distinct adhesion, see [203] and references therein), extensions to continuous heterogeneity are natural; steps in this direction have been made, for example see [88, 172]. The resulting equations, though, present a formidable mathematical and numerical challenge: for example, a doubly non-local structure, with integral terms for summation across both space and phenotype.
@@
  • IV. Discrete, continuous, or both? As noted several times, a natural application of PS-PDEs would be to describe follower-leader type behaviour, for example as can be found in invading cancer cells [243], neural crest migration [186], and wound healing [246]. While a division into followers and leaders is conceptually straightforward, this can be overly simplistic and misleading [232]. A broad spectrum of states may exist between ‘follower’ and ‘leader’, with transitions in between. PS-PDE models, therefore, can form a natural framework to describe this complexity. But questions then arise regarding the appropriateness of using continuous densities: within such systems, certain states may be confined to small subpopulations – for example, an invasion process driven by relatively few cells at the front. Turning to a fully discrete representation is one possibility, but this may also be suboptimal if other states are exhibited by vast numbers of cells. How can we form a bridge, spanning discrete to continuous populations according to phenotypic state? Some hybrid approaches offer tantalising directions, e.g. allowing individual cells to emerge from or merge with some continuous distribution, in a way that ensures mass is conserved [56].
@@
  • V. Self-organisation within phenotype-structured populations. PS-PDEs are sophisticated equations and present a formidable mathematical challenge. However, in-roads are being made: for example, we illustrated the growing arsenal of techniques to understand phenotype-structured travelling waves. Travelling waves are one well-known dynamic that can arise from PDEs, another being self-organising phenomena. [239] laid the essential groundwork here, via his well known model for morphogenesis. The now standard stability analysis relied on a near homogeneous initial state but, as noted earlier, populations are rarely homogeneous and one could expect phenotypic heterogeneity across even a superficially similar population – e.g. varying levels of signal activity. Extending pattern formation analyses to models that contain phenotype structuring is relatively straightforward under simpler binary phenotypes – e.g. see [201, 211, 181] – as the resultant PDEs are of low order and the same standard Turing-type analysis can be applied. But extending these analyses to the non-local PS-PDE framework has, to today, received little attention.
@@
  • VI. Analysis of qualitative and quantitative properties of solutions. The analysis of PS-PDEs modelling spatial spread and evolutionary dynamics of forms (7), (17), and (20) is a relatively young research field, wherein a number of fascinating and challenging open problems have emerged. Just to mention a few problems related to aspects we have focused on in this review, neither explicit nor semi-explicit solutions are known, with the exception of the case discussed in Section 4.2.3. This limits our current level of information on the quantitative properties of the solutions to these models. Moreover, formal asymptotic results on concentration phenomena across travelling waves, like those discussed here, have still to be established rigorously. In this we anticipate major difficulties to be posed due to the lack of compactness [192, 138]. Furthermore, there are no rigorous asymptotic results, of the type presented by [71], on the derivation of free-boundary problems of Hele-Shaw type from PS-PDE models of form (17), when the mobility parameter is a function of the phenotypic state. This problem appears to be far from being closed given that, as of today, related problems remain open even for systems of PDEs with cross-diffusion terms corresponding to the case where the phenotypic state is binary [169, 74]. Solving these and related problems will entail harnessing a range of tools and techniques from across different research areas for the analysis of non-linear and non-local PDEs, thus promoting cross-fertilisation of these areas. Further, it may lead to the development of new mathematical methods that could also be transferable to cognate research fields.
@@
  • VII. PS-PDEs over networks. Migration not only allows cells to infiltrate neighbouring areas but also disperse to distant sites, as happens in the metastatic spread of cancer. The question of whether cancerous cells that originate in one organ can migrate to, adapt to, and persist within another organ is of manifest interest; related questions arise in an ecological context. In such instances, continuous descriptions of space may no longer be appropriate – the distance between two organs is not easily defined – and formulations on a network may be more appropriate. Steps in this direction have been made, by describing the dynamics of the phenotype density in each network node with a non-local PDE of form (10), and adding sink/source terms for the migration across adjacent edges. Assuming fitness functions in the form of (11),(12), analytical results extend those outlined in Section 4.2.2 and highlight that spatial movement — even in discrete space settings — may provide the substrate for polymorphism, both under constant [190, 191] or phenotype-dependent migration rates [200]. The long-term persistence (or extinction) of phenotype-structured populations in such modelling framework was investigated by [126, 8], in the context of host-pathogen interactions. Nevertheless, PS-PDEs over networks currently neglect explicit spatial dynamics within each node – with the exception of a graph-like description of the habitat in each node [36] – and the inclusion of continuous spatial structures for node-specific dynamics is an open research avenue.
@@
  • VIII. Development and analysis of numerical schemes. Relying on structuring over both phenotype and physical space, PS-PDE models necessitate discretisation over multiple dimensions to be simulated. This aspect, coupled with the presence of non-local and highly non-linear terms, often makes solving numerically the equations comprised in these models prohibitively computationally expensive, especially when high-dimensional phenotypic and physical domains are simultaneously considered. As such, a significant challenge is posed when it comes to developing efficient algorithms implementing accurate numerical schemes. As discussed in Section 4.3, approaches to date have typically invoked semi-classical ideas (method of lines, finite difference, finite volume schemes, etc), which benefit from a substantial literature and flexibility. However, whether other approaches – such as pseudospectral methods [104], discontinuous Galerkin methods [65], particle methods [12], and deep-learning algorithms [37] – could be used to generate more efficient algorithms is an intriguing possibility. Beyond the development of efficient algorithms, systematic numerical analysis of discretisation schemes for PS-PDE models of cell movement of the type considered here is generally lacking in the extant literature, which would provide valuable information on accuracy of numerical solutions; this is certainly another avenue for future research.
@@
  • IX. Bridging the gap between models and data. Recent technological advances have made it possible to access a vast amount of data on the phenotypic properties of cells, both at the single-cell level and across space (see Figure 2(c)), that could be used to validate PS-PDE models of collective cell migration. While some works already leveraged proteomics data from bulk measurements, of the type shown in panel (ii) of Figure 2(c), for parameter estimation in spatially homogeneous models of phenotypically-structured cell population dynamics [11, 51], this challenge is yet to be tackled in PS-PDEs models of cell movement. Indeed spatial data come at varying levels of resolution and may be more prone to stochastic variation compared to bulk measurement data, making the task of PDE calibration more involved. This raises the question of practical identifiability of parameters in PS-PDE models, which is bound to merge complexities found when addressing this in spatially-structured phenotypically homogeneous populations [163] and in well-mixed phenotype-structured populations [40].
@@
Ethical Statement Not applicable.
@@
Code availability Python code is available on the GitHub repository https://github.com/ChiaraVilla/LorenziEtAl2025Phenotype under GNU General Public License (https://www.gnu.org/licenses/).